# Constrained Online Convex Optimization with Polyak Feasibility Steps

### v1

@@ 1 Introduction
The OCO problem has emerged as a fundamental setting for various machine learning domains, such as stochastic optimization Cesa-Bianchi et al. (2004), non-convex optimization Cutkosky et al. (2023), and online control Agarwal et al. (2019). Furthermore, the OCO problem is directly relevant to various real-world settings, including online advertising McMahan et al. (2013), internet of things Chen and Giannakis (2018), and healthcare Tewari and Murphy (2017).
Despite the significance of the OCO problem, classical methods for OCO, such as online gradient descent (OGD) (Zinkevich, 2003) and regularized follow the leader (RFTL) (Shalev-Shwartz and Singer, 2007), require orthogonal projections (or similarly costly operations) to ensure that the chosen actions are feasible. Such operations are prohibitively expensive in many applications, particularly when the action set $\mathcal{X}$ has a complicated structure. Motivated by the high computation cost of classical methods, Mahdavi et al. (2012) introduced the problem of OCO with long-term constraints, where the player is not required to ensure that the actions are feasible, but instead aims to satisfy the constraints cumulatively.
@@
Reference Regret Constraint Guarantee
Known Strictly-
Feasible Point?
Mahdavi et al. (2012) $\mathcal{O}(T^{3/4})$ $\sum_{t=1}^{T}g(x_{t})\leq 0$ No
Jenatton et al. (2016) $\mathcal{O}(T^{2/3})$ $\sum_{t=1}^{T}g(x_{t})\leq 0$ No
Yuan and Lamperski (2018)* $\mathcal{O}(T^{2/3})$ $\sum_{t=1}^{T}g(x_{t})\leq 0$ No
Yu et al. (2017)* $\mathcal{O}(\sqrt{T})$ $\sum_{t=1}^{T}g(x_{t})\leq 0$ No
Theorem 1 $\mathcal{O}(\sqrt{T})$ $g(x_{t})\leq 0,\ \forall t\in[T]$ Yes
Corollary 1 $\mathcal{O}(\sqrt{T})$ $g(x_{t})\leq 0,\ \forall t\geq t_{0}=\mathcal{O}(\log(T))$ No
Corollary 1 $\mathcal{O}(\sqrt{T})$ $\sum_{t=1}^{T}g(x_{t})\leq 0$ No
In particular, Mahdavi et al. (2012) considered the action set to be represented by the sub-level set of a constraint function, i.e. $\mathcal{X}=\{\mathbf{x}\in\mathbb{R}^{d}:g(\mathbf{x})\leq 0\}$ for convex [^1] Since the constraint function $g:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is not assumed to be differentiable, this can setting can handle multiple constraints $g_{1},...,g_{n}$ by defining $g(\mathbf{x}):=\max_{i\in[n]}g_{i}(\mathbf{x})$. $g:\mathbb{R}^{d}\rightarrow\mathbb{R}$. In this setting, Mahdavi et al. (2012) gave an algorithm that enjoys $\mathcal{O}(T^{3/4})$ regret and ensures cumulative constraint satisfaction $\sum_{t=1}^{T}g(\mathbf{x}_{t})\leq 0$, while only observing a subgradient and function value at the played actions, i.e. $g_{t}=g(\mathbf{x}_{t}),\mathbf{s}_{t}\in\partial g(\mathbf{x}_{t})$. This algorithm uses a primal-dual approach, in which it updates sequences of actions (primal variables) and constraint penalties (dual variables) to simultaneously minimize the regret and the cumulative constraint value. Primal-dual approaches have continued to be successful for this problem, as more recent algorithms of this type have been shown to enjoy $\mathcal{O}(\sqrt{T})$ regret with the same constraint guarantees and constraint feedback Yu et al. (2017).
In this work, we take a new approach to this problem by using what we call Polyak feasibility steps, which are subgradient descent steps that are taken with respect to the constraint function and use the Polyak stepsize. This approach is motivated by the fact that the Polyak stepsize is known to be effective in unconstrained convex optimization Polyak (1969). As such, we expect it to be similarly effective when applied to the constraint function in OCO. Accordingly, we design an algorithm that alternates between gradient descent steps with respect to the cost function, and Polyak feasibility steps. This algorithm differs from primal-dual approaches in that it maintains a single sequence of iterates (versus the two sequences used by primal-dual algorithms).
@@
Our complete results are shown alongside prior work in Table 1. As presented in this table, prior work has shown cumulative constraint satisfaction $\sum_{t=1}^{T}g(\mathbf{x}_{t})\leq 0$ and regret bounds as tight as $\mathcal{O}(\sqrt{T})$ Mahdavi et al. (2012); Jenatton et al. (2016); Yuan and Lamperski (2018); Yu et al. (2017). We show stronger anytime constraint satisfaction $g(\mathbf{x}_{t})\leq 0\ \forall t\in[T]$, while also guaranteeing $\mathcal{O}(\sqrt{T})$ regret. However, unlike prior work, these guarantees require that there is a known point $\mathbf{x}\in\mathbb{R}^{d}$ that is strictly-feasible $g(\mathbf{x})<0$. Nonetheless, we show that when a strictly-feasible point is not known, then we can guarantee constraint satisfaction after $\mathcal{O}(\log(T))$ rounds, as well as cumulative constraint satisfaction.
Also, note that Table 1 only includes algorithms that access the constraint function via the constraint value and subgradient at the played actions $g_{t}=g(\mathbf{x}_{t}),\mathbf{s}_{t}\in\partial g(\mathbf{x}_{t})$. Therefore, it does not include the line of literature in constrained OCO that solves a convex optimization problem with the constraint function in each round, e.g. Yu and Neely (2020); Yi et al. (2021); Guo et al. (2022). We discuss these related works in more detail in the following section.
@@ 1.1.1 Constrained OCO
Mahdavi et al. (2012) studied OCO with a fixed convex constraint function $g$, and gave an algorithm with $\mathcal{O}(\sqrt{T})$ regret and $\mathcal{O}(T^{3/4})$ cumulative violation $\sum_{t=1}^{T}g(\mathbf{x}_{t})$ that used only first-order constraint feedback, $g_{t}=g(\mathbf{x}_{t}),\mathbf{s}_{t}\in\partial g(\mathbf{x}_{t})$. This result was then generalized to $\mathcal{O}(T^{\max(\beta,1-\beta)})$ regret and $\mathcal{O}(T^{1-\beta/2})$ cumulative violation for any $\beta\in(0,1)$ by Jenatton et al. (2016). The same bounds were shown for a stronger notion of constraint violation $\sum_{t=1}^{T}[g(x_{t})]_{+}$ by Yuan and Lamperski (2018), who also guaranteed that $\sum_{t=1}^{T}([g(x_{t})]_{+})^{2}=\mathcal{O}(T^{1-\beta})$. Finally, Yu et al. (2017) showed $\mathcal{O}(\sqrt{T})$ regret and $\mathcal{O}(\sqrt{T})$ cumulative violation in the same setting. With the additional assumption that the constraint gradient is lower bounded near the constraint boundary (Assumption 1 in Mahdavi et al. (2012)), the aforementioned results can be extended to guarantee no cumulative violation $\sum_{t=1}^{T}g(\mathbf{x}_{t})\leq 0$ as stated in Table 1. We use the same assumption to show constraint satisfaction for all rounds $g(x_{t})\leq 0\ \forall t$ provided that there is a known strictly-feasible point. Furthermore, the aforementioned works use primal-dual algorithms, which are fundamentally different from our approach of Polyak feasibility steps. In particular, our approach uses the Polyak step-size and a single sequence of iterates, while primal-dual algorithms use two iterate sequences that are linked via the cost and constraint functions.
@@
There is also a line of literature on OCO with constraints that solves an optimization problem involving the constraint function in each round, e.g. Yu and Neely (2020); Yi et al. (2021, 2022); Guo et al. (2022). This differs from our algorithm and those compared in Table 1, which only access the constraint with first-order feedback at the played actions. We note that solving an optimization problem with the constraint function in each round can introduce significant computational cost.
@@
Lastly, we point out that a related line of literature considers OCO with time-varying constraints, e.g. (Neely and Yu, 2017; Yu et al., 2017; Liakopoulos et al., 2019; Castiglioni et al., 2022; Guo et al., 2022; Kolev et al., 2023). This generalizes the fixed constraints that we consider. Nonetheless, to our knowledge, none of these works improve on the guarantees shown in Table 1 for the case of non-smooth constraints, and first-order feedback. We also point out that Kolev et al. (2023) gives an algorithm that uses first-order feedback and enjoys bounds on the violation in each round, i.e. $g(x_{t})\leq\mathcal{O}(\frac{1}{\sqrt{t}}$). However, this work requires that constraint functions are smooth (i.e. have Lipschitz gradients), making it distinct from our work and those considered in Table 1. Furthermore, Kolev et al. (2023) uses a projection on to a polytope in each round, which introduces additional computational cost.
@@ 1.1.2 Projection-free OCO
In parallel to the literature on constrained OCO, there is a line of literature that considers projection-free OCO where the feasible set is not treated as the sub-level set of a function, but rather defined as an arbitrary convex set. Hazan and Kale (2012) initiated this literature by giving an algorithm that accesses the feasible set via a linear optimization oracle (LOO) instead of using projections. This is advantageous because the LOO is often computationally cheaper than the projection. LOO-based algorithms have received a significant amount of attention, e.g. (Hazan and Kale, 2012; Chen et al., 2019; Garber and Kretzu, 2020, 2022; Wang et al., 2024; Garber and Kretzu, 2024). The state-of-the-art for LOO-based algorithms with general convex costs and general convex feasible sets is $\mathcal{O}(T^{3/4})$ regret and $1$ oracle call per a round Hazan and Kale (2012). Although our methods give smaller $\mathcal{O}(\sqrt{T})$ regret, we note that the first-order feedback that we use is generally incomparable to the LOO oracle in terms of computationally complexity. Indeed, the first-order feedback is cheaper to compute for some constraint functions, while the LOO is cheaper to compute for other constraint functions. [^2] An example of this, pointed out by Garber and Kretzu (2022), is the difference in computational complexity for first-order information and LOO for the nuclear norm ball $\mathbb{B}_{*}$ and spectral norm ball $\mathbb{B}_{2}$ in the space of matrices. Computing first-order information for $\mathbb{B}_{*}$ and $\mathbb{B}_{2}$ is at worst a full-rank SVD (which is expensive) and rank-one SVD (which is cheap), respectively. For the LOO, the opposite is true in that $\mathbb{B}_{*}$ requires a rank-one SVD and $\mathbb{B}_{2}$ a full-rank SVD.
@@
There is also a growing body of literature that uses the membership oracle (MO) or separation oracle (SO) to access the feasible set, e.g. (Levy and Krause, 2019; Garber and Kretzu, 2022; Mhammedi, 2022; Lu et al., 2023; Hu et al., 2023; Mhammedi, 2024). The MO and SO are defined as follows. Given a query point $\mathbf{x}\in\mathbb{R}^{d}$, the MO specifies whether or not $\mathbf{x}$ is in the feasible set, while the SO returns a hyperplane that separates $\mathbf{x}$ from the feasible set (if $\mathbf{x}$ is not in the feasible set). If the feasible set is the sub-level set of a constraint function $g$ (as we consider), then the MO can be constructed by checking if $g(\mathbf{x})>0$, and the SO can be constructed using the first-order information at $\mathbf{x}$, i.e. $g(\mathbf{x}),\partial g(\mathbf{x})$. [^3] Given the query point $\mathbf{x}$ and first-order information at this point $g=g(\mathbf{x}),\mathbf{s}\in\partial g(\mathbf{x})$, it follows from convexity that $\{\mathbf{y}\in\mathbb{R}^{d}:g+\mathbf{s}^{\top}(\mathbf{y}-\mathbf{x})\leq 0\}$ is a separating hyperplane w.r.t. $\mathcal{X}=\{\mathbf{x}\in\mathbb{R}^{d}:g(\mathbf{x})\leq 0\}$. We use first-order information only at the played actions, which can therefore only be used to construct a separation oracle and membership oracle at the played action. This is distinct from existing MO and SO-based algorithms, which query the SO and MO at arbitrary points (not just the played actions). Furthermore, existing MO-based and SO-based algorithms require multiple oracle calls per a round. Specifically, existing MO-based algorithms use $\mathcal{O}(d\log(T))$ oracle calls per a round Lu et al. (2023); Mhammedi (2022), and existing SO-based algorithms use $\mathcal{O}(\log(T))$ oracle calls per a round Mhammedi (2022, 2024) or $\mathcal{O}(\kappa)$ oracle calls per a round Garber and Kretzu (2022). Note that we state these bounds for unrestricted $T$ and use $\kappa$ to refer to the eccentricity of the feasible set, i.e. $\kappa=R/r$ with $r\mathbb{B}\subseteq\mathcal{X}\subseteq R\mathbb{B}$, which can be arbitrarily large. The fact that our algorithm only requires $1$ oracle call per a round can result in significant performance advantages over these methods, particularly when the constraint is costly to evaluate. Furthermore, our approach is applicable to settings where there is only local constraint information available.

### v2

@@ 1 Introduction
The OCO problem has emerged as a fundamental setting for various machine learning domains, such as stochastic optimization (Cesa-Bianchi et al., 2004), non-convex optimization (Cutkosky et al., 2023), and online control (Agarwal et al., 2019). Furthermore, the OCO problem is directly relevant to various real-world settings, including online advertising (McMahan et al., 2013), internet of things (Chen & Giannakis, 2018), and healthcare (Tewari & Murphy, 2017).
Despite the significance of the OCO problem, classical methods for OCO, such as online gradient descent (OGD) (Zinkevich, 2003) and regularized follow the leader (RFTL) (Shalev-Shwartz & Singer, 2007), require orthogonal projections (or similarly costly operations) to ensure that the chosen actions are feasible. Such operations are prohibitively expensive in many applications, particularly when the action set $\mathcal{X}$ has a complicated structure. Motivated by the high computation cost of classical methods, Mahdavi et al. (2012) introduced the problem of OCO with long-term constraints, where the player is not required to ensure that the actions are feasible, but instead aims to satisfy the constraints cumulatively.
@@
Reference Regret Constraint Guarantee Known Strictly-Feasible Point?
Mahdavi et al. (2012) $\mathcal{O}(T^{3/4})$ $\sum_{t=1}^{T}g(\mathbf{x}_{t})\leq 0$ No
Jenatton et al. (2016) $\mathcal{O}(T^{2/3})$ $\sum_{t=1}^{T}g(\mathbf{x}_{t})\leq 0$ No
Yuan & Lamperski (2018)* $\mathcal{O}(T^{2/3})$ $\sum_{t=1}^{T}g(\mathbf{x}_{t})\leq 0$ No
Yu et al. (2017)* $\mathcal{O}(\sqrt{T})$ $\sum_{t=1}^{T}g(\mathbf{x}_{t})\leq 0$ No
Corollary 1 $\mathcal{O}(\sqrt{T})$ $g(\mathbf{x}_{t})\leq 0,\ \forall t\in[T]$ Yes
Corollary 2 $\mathcal{O}(\sqrt{T})$ $g(\mathbf{x}_{t})\leq 0,\ \forall t\geq t_{0}=\mathcal{O}(\log(T))$ No
Corollary 2 $\mathcal{O}(\sqrt{T})$ $\sum_{t=1}^{T}g(\mathbf{x}_{t})\leq 0$ No
In particular, Mahdavi et al. (2012) considered the action set to be represented by the sub-level set of a constraint function, i.e. $\mathcal{X}=\{x\in\mathbb{R}^{d}:g(x)\leq 0\}$ for convex [^1] Since the constraint function $g:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is not assumed to be differentiable, this setting can handle multiple constraints $g_{1},...,g_{n}$ by defining $g(\mathbf{x}):=\max_{i\in[n]}g_{i}(\mathbf{x})$. $g:\mathbb{R}^{d}\rightarrow\mathbb{R}$. In this setting, Mahdavi et al. (2012) gave an algorithm that enjoys $\mathcal{O}(T^{3/4})$ regret and ensures cumulative constraint satisfaction $\sum_{t=1}^{T}g(\mathbf{x}_{t})\leq 0$, while only observing a subgradient and function value at the played actions, i.e. $g_{t}=g(\mathbf{x}_{t}),\mathbf{s}_{t}\in\partial g(\mathbf{x}_{t})$. This algorithm uses a primal-dual approach, in which it updates sequences of actions (primal variables) and constraint penalties (dual variables) to simultaneously minimize the regret and the cumulative constraint value. Primal-dual approaches have continued to be successful for this problem, as more recent algorithms of this type have been shown to enjoy $\mathcal{O}(\sqrt{T})$ regret with the same constraint guarantees and constraint feedback (Yu et al., 2017).
In this work, we take a new approach to this problem by using what we call Polyak feasibility steps, which are subgradient descent steps that are taken with respect to the constraint function and use the Polyak stepsize. This approach is motivated by the fact that the Polyak stepsize is known to be effective in unconstrained convex optimization (Polyak, 1969). As such, we expect it to be similarly effective when applied to the constraint function in OCO. Accordingly, we design an algorithm that alternates between gradient descent steps with respect to the cost function, and Polyak feasibility steps. This algorithm differs from primal-dual approaches in that it maintains a single sequence of iterates (versus the two sequences used by primal-dual algorithms).
@@
Our complete results are shown alongside prior work in Table 1. As presented in this table, prior work has shown cumulative constraint satisfaction $\sum_{t=1}^{T}g(\mathbf{x}_{t})\leq 0$ and regret bounds as tight as $\mathcal{O}(\sqrt{T})$ (Mahdavi et al., 2012; Jenatton et al., 2016; Yuan & Lamperski, 2018; Yu et al., 2017). We show stronger anytime constraint satisfaction $g(\mathbf{x}_{t})\leq 0\ \forall t\in[T]$, while also guaranteeing $\mathcal{O}(\sqrt{T})$ regret. However, unlike prior work, these guarantees require that there is a known point $\mathbf{x}\in\mathbb{R}^{d}$ that is strictly-feasible $g(\mathbf{x})<0$. Nonetheless, we show that when a strictly-feasible point is not known, then we can guarantee constraint satisfaction after $\mathcal{O}(\log(T))$ rounds, as well as cumulative constraint satisfaction.
Also, note that Table 1 only includes algorithms that access the constraint function via the constraint value and subgradient at the played actions $g_{t}=g(\mathbf{x}_{t}),\mathbf{s}_{t}\in\partial g(\mathbf{x}_{t})$. Therefore, it does not include the line of literature in constrained OCO that solves a convex optimization problem with the constraint function in each round, e.g. (Yu & Neely, 2020; Yi et al., 2021; Guo et al., 2022). We discuss these related works in more detail in the following section.
@@ 1.1.1 Constrained OCO
Mahdavi et al. (2012) studied OCO with a fixed convex constraint function $g$, and gave an algorithm with $\mathcal{O}(\sqrt{T})$ regret and $\mathcal{O}(T^{3/4})$ cumulative violation $\sum_{t=1}^{T}g(\mathbf{x}_{t})$ that used only first-order constraint feedback, $g_{t}=g(\mathbf{x}_{t}),\mathbf{s}_{t}\in\partial g(\mathbf{x}_{t})$. This result was then generalized to $\mathcal{O}(T^{\max(\beta,1-\beta)})$ regret and $\mathcal{O}(T^{1-\beta/2})$ cumulative violation for any $\beta\in(0,1)$ by Jenatton et al. (2016). The same bounds were shown for a stronger notion of constraint violation $\sum_{t=1}^{T}[g(x_{t})]_{+}$ by Yuan & Lamperski (2018), who also guaranteed that $\sum_{t=1}^{T}([g(x_{t})]_{+})^{2}=\mathcal{O}(T^{1-\beta})$. Finally, Yu et al. (2017) showed $\mathcal{O}(\sqrt{T})$ regret and $\mathcal{O}(\sqrt{T})$ cumulative violation in the same setting. With the additional assumption that the constraint gradient is lower bounded near the constraint boundary (Assumption 1 in Mahdavi et al. (2012)), the aforementioned results can be extended to guarantee no cumulative violation $\sum_{t=1}^{T}g(\mathbf{x}_{t})\leq 0$ as stated in Table 1. We use the same assumption to show constraint satisfaction for all rounds $g(x_{t})\leq 0\ \forall t$ provided that there is a known strictly-feasible point. Furthermore, the aforementioned works use primal-dual algorithms, which are fundamentally different from our approach of Polyak feasibility steps. In particular, our approach uses the Polyak step-size and a single sequence of iterates, while primal-dual algorithms use two iterate sequences that are linked via the cost and constraint functions.
@@
There is also a line of literature on OCO with constraints that solves an optimization problem involving the constraint function in each round, e.g. (Yu & Neely, 2020; Yi et al., 2021, 2022; Guo et al., 2022). This differs from our algorithm and those compared in Table 1, which only access the constraint with first-order feedback at the played actions. We note that solving an optimization problem with the constraint function in each round can introduce significant computational cost.
@@
Lastly, we point out that a related line of literature considers OCO with time-varying constraints, e.g. (Neely & Yu, 2017; Yu et al., 2017; Liakopoulos et al., 2019; Castiglioni et al., 2022; Guo et al., 2022; Kolev et al., 2023). This generalizes the fixed constraints that we consider. Nonetheless, to our knowledge, none of these works improve on the guarantees shown in Table 1 for the case of non-smooth constraints, and first-order feedback. We also point out that Kolev et al. (2023) gives an algorithm that uses first-order feedback and enjoys bounds on the violation in each round, i.e. $g(x_{t})\leq\mathcal{O}(\frac{1}{\sqrt{t}}$). However, this work requires that constraint functions are smooth (i.e. have Lipschitz gradients), making it distinct from our work and those considered in Table 1. Furthermore, Kolev et al. (2023) uses a projection on to a polytope in each round, which introduces additional computational cost.
@@ 1.1.2 Projection-free OCO
In parallel to the literature on constrained OCO, there is a line of literature that considers projection-free OCO where the feasible set is not treated as the sub-level set of a function, but rather defined as an arbitrary convex set. Hazan & Kale (2012) initiated this literature by giving an algorithm that accesses the feasible set via a linear optimization oracle (LOO) instead of using projections. This is advantageous because the LOO is often computationally cheaper than the projection. LOO-based algorithms have received a significant amount of attention, e.g. (Hazan & Kale, 2012; Chen et al., 2019; Garber & Kretzu, 2020, 2022; Wang et al., 2024; Garber & Kretzu, 2024). The state-of-the-art for LOO-based algorithms with general convex costs and general convex feasible sets is $\mathcal{O}(T^{3/4})$ regret and $1$ oracle call per a round (Hazan & Kale, 2012). Although our methods give smaller $\mathcal{O}(\sqrt{T})$ regret, we note that the first-order feedback that we use is generally incomparable to the LOO oracle in terms of computationally complexity. Indeed, the first-order feedback is cheaper to compute for some constraint functions, while the LOO is cheaper to compute for other constraint functions. [^2] An example of this, pointed out by Garber & Kretzu (2022), is the difference in computational complexity for first-order information and LOO for the nuclear norm ball $\mathbb{B}_{*}$ and spectral norm ball $\mathbb{B}_{2}$ in the space of matrices. Computing first-order information for $\mathbb{B}_{*}$ and $\mathbb{B}_{2}$ is at worst a full-rank SVD (which is expensive) and rank-one SVD (which is cheap), respectively. For the LOO, the opposite is true in the $\mathbb{B}_{*}$ requires a rank-one SVD and $\mathbb{B}_{2}$ a full-rank SVD.
@@
There is also a growing body of literature that uses the membership oracle (MO) or separation oracle (SO) to access the feasible set, e.g. (Levy & Krause, 2019; Garber & Kretzu, 2022; Mhammedi, 2022; Lu et al., 2023; Hu et al., 2023; Mhammedi, 2024). The MO and SO are defined as follows. Given a query point $\mathbf{x}\in\mathbb{R}^{d}$, the MO specifies whether or not $\mathbf{x}$ is in the feasible set, while the SO returns a hyperplane that separates $\mathbf{x}$ from the feasible set (if $\mathbf{x}$ is not in the feasible set). If the feasible set is the sub-level set of a constraint function $g$ (as we consider), then the MO can be constructed by checking if $g(\mathbf{x})>0$, and the SO can be constructed using the first-order information at $\mathbf{x}$, i.e. $g(\mathbf{x}),\partial g(\mathbf{x})$. [^3] Given the query point $\mathbf{x}$ and first-order information at this point $g=g(\mathbf{x}),\mathbf{s}\in\partial g(\mathbf{x})$, it follows from convexity that $\{\mathbf{y}\in\mathbb{R}^{d}:g+\mathbf{s}^{\top}(\mathbf{y}-\mathbf{x})\leq 0\}$ is a separating hyperplane w.r.t. $\mathcal{X}=\{\mathbf{x}\in\mathbb{R}^{d}:g(\mathbf{x})\leq 0\}$. We use first-order information only at the played actions, which can therefore only be used to construct a separation oracle and membership oracle at the played action. This is distinct from existing MO and SO-based algorithms, which query the SO and MO at arbitrary points (not just the played actions). Furthermore, existing MO-based and SO-based algorithms require multiple oracle calls per a round. Specifically, existing MO-based algorithms use $\mathcal{O}(d\log(T))$ oracle calls per a round (Lu et al., 2023; Mhammedi, 2022), and existing SO-based algorithms use $\mathcal{O}(\log(T))$ oracle calls per a round (Mhammedi, 2022, 2024) or $\mathcal{O}(\kappa)$ oracle calls per a round (Garber & Kretzu, 2022). Note that we state these bounds for unrestricted $T$ and use $\kappa$ to refer to the eccentricity of the feasible set, i.e. $\kappa=R/r$ with $r\mathbb{B}\subseteq\mathcal{X}\subseteq R\mathbb{B}$, which can be arbitrarily large. The fact that our algorithm only requires $1$ oracle call per a round can result in significant performance advantages over these methods, particularly when the constraint is costly to evaluate. Furthermore, our approach is applicable to settings where there is only local constraint information available.